# Quantum Internet: Technologies, Protocols, and Research Challenges

### v1

@@ 4.2 Types of quantum network
According to current consensus in the field, the quantum internet is envisioned to operate alongside the classical internet, particularly during its initial phases [59]. Therefore, the categorisation of quantum networks that collectively constitute the quantum internet parallels that of the classical internet, based on the operational area’s size. Broadly, quantum networks can be categorised into modular networks, ground-based networks, and satellite-based networks.
@@
On the other hand, ground-based networks can be further subdivided according to their operational distances similar to classical networks. The most frequently discussed type of ground-based networks in current research—and those that often have experimental test beds—are metropolitan quantum networks. Metropolitan quantum networks operate over metro-scale distances [62, 63], facilitating regional connectivity within a more confined geographic area compared to their long-range counterparts.
The motivation for including satellite-based networks in the global internet stems from the fact that free space links suffer polynomial loss compared to an exponential loss in optical fibre links [64]. Therefore, satellite-based quantum networks unlock quantum communication over continental and inter-continental ranges which is otherwise challenging with purely ground-based quantum networks. However, the free space link (Forward &\&& Return link) passes through the atmosphere which induces several effects such as diffraction, absorption, and scintillation effects [65]. While research efforts are required on this front, the perseverance of entanglement in good weather conditions can be seen as a potential to include satellite-based networks in the quest for the quantum internet [66]. Given the vast distance of operation in satellite-based networks, latency also plays a huge part when considering the decoherence of quantum states. Quantum memories would play a vital role in such scenarios. Studies are ongoing on this front for the use of quantum memories in space [67].
@@ 4.3 Generations of quantum repeater
Quantum repeaters, as introduced in Section 2, are essential for the distribution of end-to-end entanglement especially in a long-distance regime. However, two major challenges complicate this task: photon loss and operational errors. Photon loss occurs when photons, the carriers of quantum information, are absorbed or scattered during transmission, preventing successful entanglement. Operational errors, on the other hand, stem from imperfections in the devices that manipulate and measure quantum states, potentially leading to incorrect results.
@@
The second generation of quantum repeaters uses HEG to correct loss errors while quantum error correction (QEC) to correct operational errors. HEG in the second generation of quantum repeaters still uses two-way signalling while QEC only requires one-way signalling which reduces the time consumption in the establishment of the end-to-end entanglement.
The third generation of quantum repeaters uses QEC to correct loss and operational errors. Subsequently, only one-way signalling is enough for the operation of the third generation of quantum repeaters which greatly reduces the time consumption in the end-to-end entanglement establishment procedure.
@@ 4.4 Routing, forwarding, and scheduling
As introduced in Section 2.1, the applications in Section 3 utilise two processes. First, transporting an arbitrary quantum state using quantum teleportation protocol. Second, swapping entanglement distributed over smaller distances with an end-to-end entanglement between two nodes or devices using an entanglement-swapping protocol. While quantum teleportation is necessary for quantum communication, establishing end-to-end entanglement unlocks most applications. It is to be noted that the two protocols are almost similar and inter-convertible with only the difference of an extra qubit and bell state measurement involved as depicted in Figure 3.
@@
For multiple end-to-end entanglement deliveries between sources-destination pairs in a quantum network, finding an optimal path using some routing metric (discussed in Section 5.1) is referred to as routing. Meanwhile, the actual execution of the entanglement swapping procedure and purification (depending upon the protocol), which considers the path found using the routing algorithm, is referred to as forwarding. The routing phase can further be categorised based on entanglement generation utilised i.e., on-demand or advanced entanglement generation (depicted in figure 9). On-demand entanglement generation calculates routing paths before initiating entanglement, whereas advanced entanglement generation bases its routes on the network topology that emerges following the probabilistic success or failure of initial link establishment. Each approach offers distinct benefits and limitations. On-demand generation allows routing decisions to be made before quantum states begin to decohere, which is advantageous. However, due to the inherently probabilistic nature of entanglement, it may require multiple attempts to secure all necessary links for a successful route. On the other hand, advanced entanglement generation relies on post-link-establishment topology data. While this method integrates more recent network information, it risks the routing decisions occurring within the decoherence window of the entangled photon pairs, which could degrade the quality of the entanglement. The choice of entanglement generation scheme significantly impacts the quantum network’s requirements. For on-demand entanglement generation, a study [80] has proposed two figures of merit. The first is the average connection time, which dictates the quantum memory requirements. The second is the average largest entanglement cluster size, which indicates the scalability of the quantum networks. In the literature, the routing phases used in on-demand and advanced entanglement generation are also referred to as proactive and reactive, respectively.
@@
Figure 9: a) Advanced entanglement generation b) On-demand entanglement generation
The issue of scheduling has recently begun to attract attention in the field of quantum networking. Scheduling involves selecting which end-to-end entangled EPR pairs to allocate to a current path request within a specific time slot while reserving EPR pairs for future path requests [81, 82, 83, 84]. This process is crucial for managing latency, particularly due to the decoherence of quantum states. Effective scheduling ensures that the quantum states are utilised efficiently, minimising the impact of decoherence on network performance and maximising the fidelity of quantum communications.
@@ 4.5 Quantum internet protocol stack
For efficient and scalable network development, a protocol stack is indispensable, as it offers a structured framework enabling independent development of each layer. Numerous research groups have put forth proposals for a quantum internet protocol stack, similar to the classical internet’s Open Systems Interconnection (OSI) model, which segregates network functions into distinct layers. In this paper, we will examine three quantum internet protocol stacks that are most frequently referenced in the scholarly literature. In Figure 10 we provide a graphical visualisation of the stacks, highlighting their differences. For more details on this specific topic, interested readers are invited to check recent relevant surveys, like [99, 100].
The initial comprehensive proposal for a layered quantum internet protocol stack dates back to 2009, introduced by Meter et al. [90, 91]. This protocol stack comprises five layers: physical entanglement, link entanglement control, error management, quantum state propagation, and application. The physical entanglement layer is tasked with generating EPR pairs between adjacent nodes, while the link entanglement control layer monitors the success and failure of these pair establishment attempts. These two layers operate across every single hop and are applied recursively to maintain connectivity through shared EPR pairs. The error management layer specifies the hops where purification is needed which ensures the high fidelity of the EPR pairs and records the outcomes. The quantum state propagation layer is responsible for establishing end-to-end entanglement by creating shared EPR pairs between the source and destination stations using the entanglement swapping procedure. This layer also communicates the results of the swapping to the end nodes, enabling them to perform the necessary single-qubit operations. Both the error management and quantum state propagation layers are similarly operated recursively across multiple hops to ensure high fidelity of the distributed entanglement. Finally, the application layer primarily manages the applications run on the quantum internet using high-fidelity EPR pairs. Recent advancements in this protocol stack have focused on ensuring synchronicity between distant nodes [92]. This study introduced a RuleSet-based quantum link bootstrapping protocol that assesses the fidelity of quantum links and their throughput. Another study implemented a quantum recursive network architecture (QRNA) alongside the RuleSet-based protocol to enhance scalability, specifically achieving multi-party entanglement and internetworking within quantum networks [93].
@@ 5.1 Simulation
First, finding a “realistic” network topology is an issue per se, which so far has been addressed inspiring by classical networks. A significant number of current research utilise the Waxman model [123] to acquire random topology of the quantum network. Others use lattices (grid and its variations), rings & spheres [124], linear chains [125], and specialised topology i.e., SURFnet [85, 126, 127] and US backbone network [73, 127].
Figure 11: The timing diagram for entanglement distribution in a centralised setting: within a period $T_{0}$, $n+1$ EPR pairs are distributed for a given set of connection requests (modified from [77])
For simplicity, simulators generally employ synchronised time slots $\tau_{n}\forall\ n\in\mathbf{N}$ which means the node clocks are assumed to be perfectly synchronised, during which the quantum states are presumed to remain coherent, as illustrated in Figure 11. Within this time slot, the distribution of end-to-end entanglement is conducted i.e., the internal and external links establishing the path between the endpoints [129, 128]. The external links are the initial EPR pairs shared between the neighbouring nodes through the quantum channels denoted by block 3 in figure 11. The internal links are within the quantum repeaters to which entanglement swapping is attempted to complete the route between the two endpoints to complete the request denoted collectively by blocks 4 & 5 in figure 11. This time slot is reiterated until a modification in the connection requests occurs, which may manifest either through the fulfilment of existing requests or the incorporation of new ones. All the operations such as quantum gate application performed are usually assumed to be executed in negligible time.
@@
As introduced in Section 4.4, routing metrics are required for the entanglement distribution. In this regard, a wide range of routing metrics has been used in the literature including throughput [128, 85, 79, 78], inverse throughput [75], continuous fidelity curves (entanglement generation fidelity vs rate) [133], end-to-end entanglement rate of the path [70], hop count [134], latency [81], E2E fidelity [135, 76, 77], and fairness [136]. The selection of routing paths using the routing metric can also be categorised using different routing algorithms such as Dijkstra [75], linear programs [137], stochastic programs [138], integer programming [139], greedy algorithms [129, 130], and AI-based routing algorithms [140, 141, 142]. Additional critical parameters may be constant or variable, and encompass link generation success rates and entanglement swapping efficiencies. The model also differentiates between two types of topology information—local [143] and global [144]—which are critical in making routing decisions for end-to-end entanglement. Furthermore, as discussed earlier the simulation incorporates two modes of heralded entanglement generation, essential for enabling entanglement swapping: on-demand [70, 128] and advanced [144, 134] as depicted in figure 9.
A majority of current research on quantum internet is done using simulation tools due to the obvious non-existent of quantum internet or networks. The notable ones are NetSquid [103], SimulaQron [104], QuNetSim [105], and SeQUeNCe [106]. A list of quantum network simulators can be found in table 6. A recent survey on the simulation tools can be found in [110].
@@ 5.2 Test beds
As discussed in Section 4, the initial simplicity of quantum internet architecture has catalysed a series of QKD network experiments worldwide. These experiments vary by several parameters including the number of nodes, the type of link used (either fibre optic or free space), the secret key generation rate, the specific QKD protocol employed, and the use of optical switching. Early implementations have typically involved short-range QKD networks and local QKD networks, employing optical components like beam splitters. These foundational networks pave the way for more complex metropolitan and backbone networks. Among these, metropolitan QKD networks have shown the most advancement in experimental implementation. A detailed listing of experimentally implemented QKD-based networks is provided in [32].
The higher stages of quantum internet that require end-to-end entanglement distribution are still in nascent stages. Experimental realisation of these types of networks is currently limited to a few nodes with varying link lengths and physical systems of qubits. A non-exhaustive list of entanglement-based quantum network test beds is provided in the table 7.
@@ 6 Standardisation
To the best of our knowledge, the only global standardisation initiative also covering more general scenarios is the Quantum Internet Research Group [^9] https://www.irtf.org/qirg.html (QIRG), established within the Internet Research Task Force (IRTF), which is the twin on the Internet Engineering Task Force (IETF) focusing on longer-term research issues. The group so far has published two information Request For Comment (RFC) documents describing possible use cases and application scenarios for the Quantum Internet [145] and laying down some high-level principles for the design of the architecture of the Quantum Internet [146]. Despite the importance of their role in clarifying the motivation and terminology in a tutorial manner, and being a source of inspiration to find new research challenges, neither of these documents provides a reference model or interfaces/protocols, which leaves the issue open.
@@ 7 Challenges and future directions
In this context, we outline both challenges and future research opportunities. Specifically, we first identify the key technological enablers that are crucial for the design and implementation of quantum networks (Section 7.1). Subsequently, we discuss future research directions that could significantly propel advancements in this field, thereby catalysing the maturation and expansion of the quantum internet (Section 7.2).
7.1 Challenges: Technology enablers
@@
Quantum repeaters, as key components in quantum networks, are critical for advancements in the field. Unfortunately, a fully functional physical quantum repeater has yet to be realised. Currently, a quantum repeater is typically an experimental station that simulates the behaviour of a quantum repeater node. The miniaturisation and advancement of quantum repeaters represent a significant challenge, closely tied to broader developments in quantum technology.
@@
7.2 Research challenges and Future directions
Realising the quantum internet is a global interdisciplinary process. Therefore, there are several open research challenges associated with the quantum internet, from many perspectives and covering heterogeneous areas of expertise.
@@ 7.2 Research challenges and Future directions
As outlined in Section 4.5, while various quantum internet protocol stacks are discussed in the literature, a consensus on the reference model has not yet been established. The lack of a universally superior physical system highlights the need for a well-defined interface between the foundational hardware and the software layers above. Observing the experimental realisation of these proposed stacks is crucial. Progress in this area includes the experimental implementation of link and physical layer protocols within a network consisting of just two nodes based on diamond NV centres [155], where entangled states were delivered with the fidelity specified by the user. Additionally, the unique properties of the quantum internet challenge the suitability of the universal layer model that has been effective in the classical internet. This ongoing debate emphasises the necessity for a customised approach tailored to the specific needs and capabilities of quantum communication technologies.
As detailed in Section 4.4, the distribution of end-to-end entanglement, essential for realising the quantum internet, presents a significant unresolved challenge. Various strategies addressing routing, forwarding, and scheduling have been investigated, yet a comprehensive understanding remains elusive. There is currently no definitive model that fully addresses all the complexities of the optimisation problem involved in quantum entanglement distribution, a problem which has no direct classical counterpart. A significant challenge within this model is the inherent non-determinism; there is a probability of failure in both the establishment of EPR pairs used for the swapping procedure and in the entanglement purification process. To address these challenges, a cost-vector formalism model has been proposed [101] recently. The essence of this model is to characterise each EPR pair with two probability weights: the transmission probability, which reflects the likelihood of successfully establishing an EPR pair, and the coherence probability, which indicates the likelihood that an EPR pair will remain coherent and thus suitable for distribution to endpoints. Clearly, further exploration in this area is necessary to more accurately model and optimise the complex processes involved in routing and distributing entanglement within quantum networks. Furthermore, the time-slot structure proposed in existing literature, which is crucial for maintaining qubit coherence, hinges on the choice of the physical system used for quantum memory. This choice is itself a significant area of research, reflecting the intricate dependencies and innovations required in the development of quantum internet technologies.
@@
Given the fragility of quantum states, entanglement purification and quantum error correction techniques are vital for addressing the challenges inherent in quantum internet. Quantum error correction, in particular, is a broad field of research that intersects significantly with the development of robust quantum computing systems. However, these techniques require additional resource consumption, which presents a trade-off. Ongoing research is focused on optimising the integration of these techniques within the network and application layer protocols of the quantum internet. This research aims to enhance the reliability and efficiency of quantum communications while managing the costs associated with these advanced error mitigation strategies.
@@
As mentioned briefly in Section 2, multi-partite entanglement, as covered by Dür et al.’s protocol stack, can be employed in quantum networks to achieve the same objectives traditionally aimed with bi-partite entanglement. This area, particularly in comparison to bi-partite entanglement, is significantly under-studied and warrants further research. A notable feature of entanglement generation using GHZ protocols is the presence of a super-critical region where the entanglement generation rate remains constant regardless of distance [158, 159]. This contrasts sharply with the exponential decay in entanglement generation rate associated with bell-state protocols, which occurs even with multi-path routing given that the bell measurement success probability $q<1$.

### v2

@@ 4.2 Types of quantum network
According to the current consensus in the field, the quantum internet is envisioned to operate alongside the classical internet, particularly during its initial phases [62]. Therefore, the categorisation of quantum networks that collectively constitute the quantum internet parallels that of the classical internet, based on the operational area’s size. Broadly, quantum networks can be categorised into modular networks, ground-based networks, and satellite-based networks.
@@
On the other hand, ground-based networks can be further subdivided according to their operational distances, similar to classical networks. The most frequently discussed type of ground-based networks in current research—and those that often have experimental test beds—are metropolitan quantum networks. Metropolitan quantum networks operate over metro-scale distances [65, 66], facilitating regional connectivity within a more confined geographic area compared to their long-range counterparts.
The motivation for including satellite-based networks in the global internet stems from the fact that free space links suffer polynomial loss compared to an exponential loss in optical fibre links [67]. Therefore, satellite-based quantum networks unlock quantum communication over continental and intercontinental ranges, which is otherwise challenging with purely ground-based quantum networks. However, the free space link (Forward &\&& Return link) passes through the atmosphere, which induces several effects, such as diffraction, absorption, and scintillation effects [68]. While research efforts are required on this front, the perseverance of entanglement in good weather conditions can be seen as a potential to include satellite-based networks in the quest for the quantum internet [69]. Given the vast distance of operation in satellite-based networks, latency also plays a huge part when considering the decoherence of quantum states. Quantum memories would play a vital role in such scenarios. Studies are ongoing on this front for the use of quantum memories in space [70].
@@ 4.3 Generations of quantum repeater
Quantum repeaters, as introduced in Section 2, are essential for the distribution of end-to-end entanglement, especially in a long-distance regime. However, two major challenges complicate this task: photon loss and operational errors. Photon loss occurs when photons, the carriers of quantum information, are absorbed or scattered during transmission, preventing successful entanglement. Operational errors, on the other hand, stem from imperfections in the devices that manipulate and measure quantum states, potentially leading to incorrect results.
@@
The second generation of quantum repeaters uses HEG to correct loss errors while quantum error correction (QEC) to correct operational errors. HEG in the second generation of quantum repeaters still uses two-way signalling, while QEC only requires one-way signalling, which reduces the time consumption in the establishment of the end-to-end entanglement.
The third generation of quantum repeaters uses QEC to correct loss and operational errors. Subsequently, only one-way signalling is enough for the operation of the third generation of quantum repeaters, which greatly reduces the time consumption in the end-to-end entanglement establishment procedure.
@@ 4.4 Routing, forwarding, and scheduling
As introduced in Section 2.1, the applications in Section 3 utilise two processes. First, transporting an arbitrary quantum state using a quantum teleportation protocol. Second, swapping entanglement is distributed over smaller distances with an end-to-end entanglement between two nodes or devices using an entanglement-swapping protocol. While quantum teleportation is necessary for quantum communication, establishing end-to-end entanglement unlocks most applications. It is to be noted that the two protocols are almost similar and inter-convertible with only the difference of an extra qubit and bell state measurement involved as depicted in Figure 3.
@@
For multiple end-to-end entanglement deliveries between source-destination pairs in a quantum network, finding an optimal path using some routing metric (discussed in Section 5.1) is referred to as routing.
Meanwhile, the actual execution of the entanglement swapping procedure and purification (depending upon the protocol), which considers the path found using the routing algorithm, is referred to as forwarding. The routing phase can further be categorised based on entanglement generation utilised, that is, on-demand or advanced entanglement generation (depicted in figure 9). On-demand entanglement generation calculates routing paths before initiating entanglement, whereas advanced entanglement generation bases its routes on the network topology that emerges following the probabilistic success or failure of initial link establishment. Each approach offers distinct benefits and limitations. On-demand generation allows routing decisions to be made before quantum states begin to decohere, which is advantageous. However, due to the inherently probabilistic nature of entanglement, it may require multiple attempts to secure all necessary links for a successful route. On the other hand, advanced entanglement generation relies on post-link-establishment topology data. While this method integrates more recent network information, it risks the routing decisions occurring within the decoherence window of the entangled photon pairs, which could degrade the quality of the entanglement. The choice of entanglement generation scheme significantly impacts the quantum network’s requirements. For on-demand entanglement generation, a study [83] has proposed two figures of merit. The first is the average connection time, which dictates the quantum memory requirements. The second is the average largest entanglement cluster size, which indicates the scalability of the quantum networks. In the literature, the routing phases used in on-demand and advanced entanglement generation are also referred to as proactive and reactive, respectively.
@@
Figure 9: a) Advanced entanglement generation: The Path is chosen out of a reduced topology constructed out of successful entanglement links. b) On-demand entanglement generation: Successful entanglement links are established on a chosen path.
The issue of scheduling has recently begun to attract attention in the field of quantum networking. Scheduling involves selecting which end-to-end entangled EPR pairs to allocate to a current path request within a specific time-slot (discussed in Section 5.1) while reserving EPR pairs for future path requests [84, 85, 86, 87]. This process is crucial for managing latency, particularly due to the decoherence of quantum states. Effective scheduling ensures that the quantum states are utilised efficiently, minimising the impact of decoherence on network performance and maximising the fidelity of quantum communications.
@@ 4.5 Quantum internet protocol stack
For efficient and scalable network development, a protocol stack is indispensable, as it offers a structured framework enabling the independent development of each layer. Numerous research groups have put forth proposals for a quantum internet protocol stack, similar to the classical internet’s Open Systems Interconnection (OSI) model, which segregates network functions into distinct layers. In this paper, we will examine three quantum internet protocol stacks that are most frequently referenced in the scholarly literature. In Figure 10, we provide a graphical visualisation of the stacks, highlighting their differences. For more details on this specific topic, interested readers are invited to check recent relevant surveys, like [102, 103].
The initial comprehensive proposal for a layered quantum internet protocol stack dates back to 2009, introduced by Meter et al. [93, 94]. This protocol stack comprises five layers: physical entanglement, link entanglement control, error management, quantum state propagation, and application. The physical entanglement layer is tasked with generating EPR pairs between adjacent nodes, while the link entanglement control layer monitors the success and failure of these pair establishment attempts. These two layers operate across every single hop and are applied recursively to maintain connectivity through shared EPR pairs. The error management layer specifies the hops where purification is needed, which ensures the high fidelity of the EPR pairs and records the outcomes. The quantum state propagation layer is responsible for establishing end-to-end entanglement by creating shared EPR pairs between the source and destination stations using the entanglement swapping procedure. This layer also communicates the results of the swapping to the end nodes, enabling them to perform the necessary single-qubit operations. Both the error management and quantum state propagation layers are similarly operated recursively across multiple hops to ensure high fidelity of the distributed entanglement. Finally, the application layer primarily manages the applications run on the quantum internet using high-fidelity EPR pairs. Recent advancements in this protocol stack have focused on ensuring synchronicity between distant nodes [95]. This study introduced a RuleSet-based quantum link bootstrapping protocol that assesses the fidelity of quantum links and their throughput. Another study implemented a quantum recursive network architecture (QRNA) alongside the RuleSet-based protocol to enhance scalability, specifically achieving multi-party entanglement and internetworking within quantum networks [96].
@@ 5.1 Simulation
First, finding a “realistic” network topology is an issue per se, which so far has been addressed inspiring by classical networks.
A significant amount of current research utilises the Waxman model [124] to acquire the random topology of the quantum network. Others use lattices (grid and its variations) [79, 80], rings & spheres [125], linear chains [126], and specialised topology, that is, SURFnet [88, 127, 128] and US backbone network [76, 128].
Figure 11: Timing diagram for centralized entanglement distribution. In each period $T_{0}$, a total of $n+1$ EPR pairs are allocated for a given set of connection requests, with the $i^{th}$ pair distributed in its designated time-slot $\tau_{i}$. Routes are determined at the start of the period during the $\tau_{r}$ slot and remain fixed throughout. A new period is initiated whenever a change in the connection requests occurs (modified from [80])
For simplicity, simulators generally employ synchronised time slots $\tau_{n}\forall\ n\in\mathbf{N}$, which means the node clocks are assumed to be perfectly synchronised, during which the quantum states are presumed to remain coherent, as illustrated in Figure 11. Within this time slot, the distribution of end-to-end entanglement is conducted, that is, the internal and external links establishing the path between the endpoints [130, 129]. The external links are the initial EPR pairs shared between the neighbouring nodes through the quantum channels denoted by block 3 in figure 11. The internal links are within the quantum repeaters to which entanglement swapping is attempted to complete the route between the two endpoints to complete the request denoted collectively by blocks 4 & 5 in figure 11. This time slot is reiterated until a modification in the connection requests occurs, which may manifest either through the fulfilment of existing requests or the incorporation of new ones [80]. All the operations such as quantum gate application performed are usually assumed to be executed in negligible time.
@@
As introduced in Section 4.4, routing metrics are required for entanglement distribution. In this regard, a wide range of routing metrics has been used in the literature including throughput [129, 88, 82, 81], inverse throughput [78], continuous fidelity curves (entanglement generation fidelity vs rate) [134], end-to-end entanglement rate of the path [73], hop count [135], latency [84], E2E fidelity [136, 79, 80], and fairness [137]. The selection of routing paths using the routing metric can also be categorised using different routing algorithms such as Dijkstra-based [78, 79, 80], linear programs [138], stochastic programs [139], integer programming [140], greedy algorithms [130, 131], and AI-based routing algorithms [141, 142, 143]. Additional critical parameters may be constant or variable and encompass link generation success rates and entanglement swapping efficiencies. The model also differentiates between two types of topology information—local [144] and global [145]—which are critical in making routing decisions for end-to-end entanglement. Furthermore, as discussed earlier, the simulation incorporates two modes of heralded entanglement generation, essential for enabling entanglement swapping: on-demand [73, 129] and advanced [145, 135], as depicted in figure 9.
A majority of current research on the quantum internet is done using simulation tools due to the obvious non-existence of the quantum internet or networks. The notable ones are NetSquid [106], SimulaQron [107], QuNetSim [108], and SeQUeNCe [109]. A list of quantum network simulators can be found in table 6. A recent survey on the simulation tools can be found in [113].
@@ 5.2 Test beds
As discussed in Section 4, the initial simplicity of quantum internet architecture has catalysed a series of QKD network experiments worldwide. These experiments vary by several parameters, including the number of nodes, the type of link used (either fibre optic or free space), the secret key generation rate, the specific QKD protocol employed, and the use of optical switching. Early implementations have typically involved short-range QKD networks and local QKD networks, employing optical components like beam splitters. These foundational networks pave the way for more complex metropolitan and backbone networks. Among these, metropolitan QKD networks have shown the most advancement in experimental implementation. A detailed listing of experimentally implemented QKD-based networks is provided in [35].
The higher stages of quantum internet that require end-to-end entanglement distribution are still in nascent stages. The experimental realisation of these types of networks is currently limited to a few nodes with varying link lengths and physical systems of qubits. A non-exhaustive list of entanglement-based quantum network test beds is provided in the table 7.
@@ 6 Standardisation
To the best of our knowledge, the only global standardisation initiative also covering more general scenarios is the Quantum Internet Research Group [^10] https://www.irtf.org/qirg.html (QIRG), established within the Internet Research Task Force (IRTF), which is the twin on the Internet Engineering Task Force (IETF) focusing on longer-term research issues. The group so far has published two information Request For Comment (RFC) documents describing possible use cases and application scenarios for the Quantum Internet [146] and laying down some high-level principles for the design of the architecture of the Quantum Internet [147]. Despite the importance of their role in clarifying the motivation and terminology in a tutorial manner and being a source of inspiration to find new research challenges, neither of these documents provides a reference model or interfaces/protocols, which leaves the issue open.
@@ 7 Challenges and future directions
In this context, we outline both challenges and future research opportunities. Specifically, we first identify the key technological enablers that are crucial for the design and implementation of quantum networks (Section 7.1). Subsequently, we discuss future research directions that could significantly propel advancements in this field, thereby catalysing the maturation and expansion of the quantum internet (Section 7.2). Finally, we explore the potential for quantum and classical networks to coexist and discuss the interplay between these systems in achieving harmonious integration (Section 7.3).
7.1 Technology enablers
@@
Quantum repeaters, as key components in quantum networks, are critical for advancements in the field.
Unfortunately, a fully functional physical quantum repeater has yet to be realised. Currently, a quantum repeater is typically an experimental station that simulates the behaviour of a quantum repeater node. The miniaturisation and advancement of quantum repeaters represent a significant challenge, closely tied to broader developments in quantum technology.
@@
7.2 Research challenges
Realising the quantum internet is a global interdisciplinary process. Therefore, there are several open research challenges associated with the quantum internet from many perspectives and covering heterogeneous areas of expertise.
@@ 7.2 Research challenges
As outlined in Section 4.5, while various quantum internet protocol stacks are discussed in the literature, a consensus on the reference model has not yet been established. The lack of a universally superior physical system highlights the need for a well-defined interface between the foundational hardware and the software layers above. Observing the experimental realisation of these proposed stacks is crucial. Progress in this area includes the experimental implementation of link and physical layer protocols within a network consisting of just two nodes based on diamond NV centres [156], where entangled states were delivered with the fidelity specified by the user.
Additionally, the unique properties of the quantum internet challenge the suitability of the universal layer model that has been effective in the classical internet. This ongoing debate emphasises the necessity for a customised approach tailored to the specific needs and capabilities of quantum communication technologies.
As detailed in Section 4.4, the distribution of end-to-end entanglement, essential for realising the quantum internet, presents a significant unresolved challenge. Various strategies addressing routing, forwarding, and scheduling have been investigated, yet a comprehensive understanding remains elusive. There is currently no definitive model that fully addresses all the complexities of the optimisation problem involved in quantum entanglement distribution, a problem which has no direct classical counterpart. A significant challenge within this model is the inherent non-determinism; there is a probability of failure in both the establishment of EPR pairs used for the swapping procedure and in the entanglement purification process. To address these challenges, a cost-vector formalism model has been proposed [104] recently. The essence of this model is to characterise each EPR pair with two probability weights: the transmission probability, which reflects the likelihood of successfully establishing an EPR pair, and the coherence probability, which indicates the likelihood that an EPR pair will remain coherent and thus suitable for distribution to endpoints. Further exploration in this area is necessary to more accurately model and optimise the complex processes involved in routing and distributing entanglement within quantum networks. Furthermore, the time-slot structure proposed in existing literature, which is crucial for maintaining qubit coherence, hinges on the choice of the physical system used for quantum memory. This choice is itself a significant area of research, reflecting the intricate dependencies and innovations required in the development of quantum internet technologies.
@@
Given the fragility of quantum states, entanglement purification and quantum error correction techniques are vital for addressing the challenges inherent in the quantum internet [159, 160]. Quantum error correction, in particular, is a broad field of research that intersects significantly with the development of robust quantum computing systems. However, these techniques require additional resource consumption, which presents a trade-off. Ongoing research is focused on optimising the integration of these techniques within the network and application layer protocols of the quantum internet. This research aims to enhance the reliability and efficiency of quantum communications while managing the costs associated with these advanced error mitigation strategies.
@@
As mentioned briefly in Section 2, multi-partite entanglement, as covered by Dür et al.’s protocol stack, can be employed in quantum networks to achieve the same objectives traditionally aimed with bi-partite entanglement. This area, particularly in comparison to bi-partite entanglement, is significantly understudied and warrants further research [161, 162, 163]. A notable feature of entanglement generation using GHZ protocols is the presence of a supercritical region where the entanglement generation rate remains constant regardless of distance [164, 165]. This contrasts sharply with the exponential decay in entanglement generation rate associated with bell-state protocols, which occurs even with multi-path routing given that the bell measurement success probability $q<1$.
7.3 Interplay with classical networks
Given the inherent differences in how quantum and classical networks operate, achieving their coexistence on a shared physical infrastructure—as outlined in Section 1—poses a significant challenge. One immediate issue is the simultaneous transmission of quantum and classical signals over the same optical fibers. While classical systems rely on high-intensity signals supported by amplifiers and multiplexing schemes, quantum signals—often at the single-photon level—are extremely weak and highly susceptible to noise such as Raman scattering from adjacent classical channels [166]. To protect quantum channels from such interference without compromising classical performance, advanced optical filtering and the development of low-noise components are essential. Moreover, quantum protocols demand synchronization with precisions down to the nanosecond or even picosecond level—requirements that far exceed those of typical classical networks [167]. Experimental setups have adopted protocols like the White Rabbit system to achieve the necessary synchronization [168], yet integrating these high-precision timing solutions into existing classical infrastructures remains a formidable task. Additionally, quantum networks inherently utilize two parallel channels: a quantum channel for transmitting qubits and a classical channel for control, coordination, and supplementary data. Integrating these channels requires managing a dual protocol stack, where the classical control plane must orchestrate complex quantum operations such as entanglement generation and swapping [167, 169]. This duality further increases network management complexity, necessitating the development of novel control algorithms and standardized interfaces for efficient resource allocation, scheduling, and routing across both domains.